What is the energy of a 10kg mass position at 100m above the surface of the earth

1
Expert's answer
2022-01-05T08:16:19-0500
"E=GPE+KE=mgh+0=mgh,""E=10\ kg\times9.8\ \dfrac{m}{s^2}\times100\ m=9800\ J."